OpenStudy (anonymous):

Using the completing-the-square method, find the vertex of the function f(x) = 2x2 − 8x + 6 and indicate whether it is a minimum or a maximum and at what point. Maximum at (2, –2) Minimum at (2, –2) Maximum at (2, 6) Minimum at (2, 6)

Mehek (mehek14):

since the first term is positive = \(2x^2\) that means it is opening upwards so the vertex would be the minimum that eliminates A and C

Mehek (mehek14):

to find the vertex, use \(\dfrac{-b}{2a}\) in your equation, you have b = -8 a = 2 \(\dfrac{-(-8)}{2*2}=\dfrac{8}{4}=2\) so x = 2

Mehek (mehek14):

plug x = 2 into the equation

Mehek (mehek14):

@boots_2000 can you do that?

OpenStudy (anonymous):

Yeah so is it maximum 2,6?

OpenStudy (anonymous):

@Mehek14

OpenStudy (irishboy123):

it says "Using the completing-the-square method"

Mehek (mehek14):

no did you plug in x = 2 in the equation

OpenStudy (anonymous):

yeah

OpenStudy (welshfella):

using completing the square:- = 2(x^2 - 4x + 3) = 2[(x - 2)^2 - c + 3] can you tell me the value of c - can you remember from the last post?

Mehek (mehek14):

\(f(2)=2*2^2-8*2+6\\2^2=4\\2*4=8\\8*2=16\\8-16=-8\) add \(-8+6\)

OpenStudy (irishboy123):

+1 @welshfella

OpenStudy (anonymous):

so its minimum? lol

Mehek (mehek14):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Mehek14 since the first term is positive = \(2x^2\) that means it is opening upwards so the vertex would be the minimum that eliminates A and C \(\color{#0cbb34}{\text{End of Quote}}\)

OpenStudy (welshfella):

its a minimum but what are the coordinates at the minimum?

OpenStudy (anonymous):

2,6

Mehek (mehek14):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Mehek14 \(f(2)=2*2^2-8*2+6\\2^2=4\\2*4=8\\8*2=16\\8-16=-8\) add \(-8+6\) \(\color{#0cbb34}{\text{End of Quote}}\)

OpenStudy (welshfella):

No mehek has worked the y coordinate for you you can also get from the completing the square method

OpenStudy (anonymous):

so 2,-2

Mehek (mehek14):

yes

OpenStudy (anonymous):

could you guys come and help me quick when you're all done here?

OpenStudy (welshfella):

I'll carry from where i left off 2[(x - 2)^2 - c + 3] c = 4 because -2^2 = + 4 = 2(x - 2)^2 - 1) = 2(x - 2)^2 - 2 so the vertex is at ( 2,-2) you get this by comparing your expression with the genarl form for the vertex

OpenStudy (anonymous):

please

OpenStudy (welshfella):